OpenStudy (anonymous):

Which of the following values of k makes the polynomial a perfect square? x2 + 11x + k a. 2/121 b.121 c.121/4 d.121/2

OpenStudy (callisto):

Perfect square: \((a+b)^2 = a^2+2ab+b^2\) Now, you have x^2 + 11x + k k is b^2 in the general form and a = x. Consider the middle term, 2ab = 11x -(1) Put a=x and \(b=\sqrt{k}\) into (1) 2(x)(\(\sqrt{k}\)) = 11x \(2\sqrt{k}\)=11 Now, solve k and get your answer!

OpenStudy (anonymous):

Im so confused..

OpenStudy (anonymous):

I need to have this done by 4:30 & im no where near done

OpenStudy (callisto):

Which part are you confused at?

OpenStudy (anonymous):

on everything.

OpenStudy (callisto):

Do you know the identity of perfect square?

OpenStudy (anonymous):

no not at all. Im returning my friend a favor by doing his work & im trying to impress him

OpenStudy (callisto):

But you are like.... asking us to do some work for you to impress him.... Perfect square expansion: \((a\pm b)^2 = a^2 \pm 2ab + b^2\) You can get this identity by expanding the left side. Try to expand (a+b)(a+b). What do you get?

OpenStudy (jiteshmeghwal9):

\[(x+k)^2=x^2+2(k)(x)+k^2\]\[x^2+2kx+k^2=x^2+11x+k\]

OpenStudy (anonymous):

never mind i found the asnwer

OpenStudy (jiteshmeghwal9):